OpenStudy (anonymous):

If f(x) = x + 5, what does f(8) represent?

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

The value of x when (x + 5) = 8 The value of 8f(x) when x = 8 The value of f(x) when (x + 5) = 8 The value of (x + 5) when x = 8

OpenStudy (phi):

If f(x) = x + 5, what does f(8) represent? remember f(x) is the "short name" for the formula x+5 (it's not very much shorter, but for hairy, complicated formulas it is very helpful) f(8) is short for: get the formula (the x+5), and everywhere you see x in the formula, put in 8 instead of x in other words, f(8) means the same thing as 8+5

OpenStudy (phi):

can you figure out the answer ?

OpenStudy (anonymous):

Is it A?

OpenStudy (phi):

f(8) means replace x with 8. another way to say that is x=8 in the formula

OpenStudy (phi):

the correct answer will have x=8 in it.

OpenStudy (anonymous):

Ohh, it's D, right? Thanks soo much! I appreciate it! :)

OpenStudy (phi):

yes D.

OpenStudy (anonymous):

Thanks soo much! :D <3